OpenStudy (anonymous):

In the following equation, identify the x-intercepts in the graph y = 3x^2 + 19x – 40

OpenStudy (lgbasallote):

factor it out

OpenStudy (anonymous):

i got (3x+8)(x-5) 3x+8=0 x-5=0 x=8/3 x=5 is that right?

jimthompson5910 (jim_thompson5910):

let's check to see if (3x+8)(x-5) is the correct factorization (3x+8)(x-5) 3x(x-5)+8(x-5) 3x^2-15x+8x-40 3x^2-7x-40 Since the result above is NOT 3x^2 + 19x – 40 (which is the original expression), this means that 3x^2 + 19x – 40 does NOT factor to (3x+8)(x-5)

OpenStudy (anonymous):

oh yeahh. i cant figure it outt

jimthompson5910 (jim_thompson5910):

The most direct way (without guessing and checking) is to use the quadratic formula

jimthompson5910 (jim_thompson5910):

Are you familiar with it?

OpenStudy (anonymous):

ehh not off the top of my head

jimthompson5910 (jim_thompson5910):

The quadratic formula is \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\]

OpenStudy (anonymous):

oh okay ik what it is

jimthompson5910 (jim_thompson5910):

In the case of \[\Large y = 3x^2 + 19x – 40\] a = 3, b = 19 and c = -40

OpenStudy (anonymous):

i got it! thank youu!!

jimthompson5910 (jim_thompson5910):

sure thing

jimthompson5910 (jim_thompson5910):

tell me what you get this time

OpenStudy (anonymous):

5/3 and 8

jimthompson5910 (jim_thompson5910):

close

jimthompson5910 (jim_thompson5910):

one of those solutions is off by a sign

OpenStudy (anonymous):

-5/3?

jimthompson5910 (jim_thompson5910):

no, it's the other one actually, the 8 should be -8

OpenStudy (anonymous):

ohhh lol thankss!

jimthompson5910 (jim_thompson5910):

here's why \[\Large x = \frac{-b\pm\sqrt{b^2-4ac}}{2a}\] \[\Large x = \frac{-(19)\pm\sqrt{(19)^2-4(3)(-40)}}{2(3)}\] \[\Large x = \frac{-19\pm\sqrt{361-(-480)}}{6}\] \[\Large x = \frac{-19\pm\sqrt{841}}{6}\] \[\Large x = \frac{-19+\sqrt{841}}{6} \ \text{or} \ x = \frac{-19-\sqrt{841}}{6}\] \[\Large x = \frac{-19+29}{6} \ \text{or} \ x = \frac{-19-29}{6}\] \[\Large x = \frac{10}{6} \ \text{or} \ x = \frac{-48}{6}\] \[\Large x = \frac{5}{3} \ \text{or} \ x = -8\] \[\Large x = -8 \ \text{or} \ x = \frac{5}{3}\]

OpenStudy (anonymous):

i gotcha!

jimthompson5910 (jim_thompson5910):

alright great